Interarrival time of mails are exponentially with a mean 10 minutes what is the probability 5 mails arrive in 20 minutes?


1
Expert's answer
2021-06-07T18:43:54-0400

Given "\lambda=1/10=0.1"


"P(x\leq\dfrac{20}{5})=1-e^{0.1(4)}\approx0.3297"

The probability 5 mails arrive in 20 minutesn is "0.3297."